Need help how to solve f(x)=(x-1)^2;g(x)=3-x
Answer by stanbon(75887)

f(x)=(x-1)^2;g(x)=3-x
--------------------
(x-1)^2 = 3-x
x^2 -2x + 1 = 3-x
x^2 -x -2 = 0
Factor:
(x-2)(x+1) = 0
-----
x = 2 or x = -1
====================
Cheers,
Stan H.
=============